### 19 Feb 2024 v2

# Fusing Individualized Treatment Rules Using Secondary Outcomes: Supplementary Materials

Fusing Individualized Treatment Rules Using Secondary Outcomes:
Supplementary Materials
In the supplementary material, we provide additional implementation details and experiment results in the simulation study and real data analysis in Section A. In addition, we provide additional theoretical results and their proof in Section B.
Appendix A ADDITIONAL DETAILS AND RESULTS FOR SECTIONS LABEL:sec:simulation AND LABEL:sec:realdata
In this section, we provide some details about implementation and experiment results in the simulation study of Section LABEL:sec:simulation and in the real data analysis of Section LABEL:sec:realdata.
A.1 Implementation Details
To find the optimal solutions for FITR-IntL and FITR-Ramp, we use the scipy package in Python. Our experiments show that the Powell algorithm generally produces better optimization results than DCA for FITR-Ramp, so we use the function minimize(method=‘Powell’) to minimize the loss function. For FITR-IntL, we use the function minimize(method=‘BFGS’) with closed-form gradients specified for optimization.
The tuning parameter $\lambda_{1n}$ is first chosen with cross-validation when estimating the ITR using SepL for each reward $R_{1},\dots,R_{K}$. Subsequently, the parameter $\mu_{1n}$ in FITR-IntL or the parameters $\mu_{1n}$ and $\kappa_{1n}$ in FITR-Ramp are tuned simultaneously with cross-validation while $\lambda_{1n}$ is kept fixed. The parameter $\sigma_{1n}$ in the Gaussian kernel is chosen as the median of the distances between all covariate pairs, which is a common heuristic when the sample size is not too large (Garreau et al.,, 2017).
A.2 Additional Details for Data Generation and Value Function Estimation in Section LABEL:sec:simulation
The first two covariates are important variables that affect the outcomes and are generated as $X_{ij}\overset{i.i.d.}{\sim}Unif(-1,1)$ for all $i=1,\dots,n$ and $j=1,2$. Let $X_{i3}=0.8X_{i3}^{\prime}+X_{i1}$, where $X_{i3}^{\prime}\overset{i.i.d.}{\sim}Unif(-1,1)$ so that $X_{i3}$ is correlated with $X_{i1}$. The remaining covariates are independently generated as $X_{ij}\overset{i.i.d.}{\sim}Unif(-1,1)$ for all $j=4,\dots,d$. Each patient’s noise variable $(\epsilon_{1},\dots,\epsilon_{K})$ follows a mean-zero multivariate normal distribution, where the covariance matrix has 0.2 on its diagonal and 0.1 on its off-diagonal entries.
The value functions are numerically calculated from an independent test set of size 100,000. The disagreement rate between two decision functions $\widehat{f}_{1n}$ and $f^{*}_{k}$ is estimated by averaging ${\mathds{1}}\{\widehat{f}_{1n}f^{*}_{k}<0\}$ on this test set. The optimal value $\mathcal{V}_{k}^{*}$ is obtained by averaging the rewards of all patients in the test set when the treatment is taken as $\operatorname*{arg\,max}_{A}T_{k}(\bm{X},A)$ for the $k$th outcome. The value function of each estimated ITR is obtained by following the corresponding treatment rule. The misclassification rate of $\widehat{f}_{1n}$ is estimated by averaging ${\mathds{1}}\{\widehat{f}_{1n}f^{*}_{1}<0\}$ on the test set.
A.3 Additional Results for Estimating FITR in Section LABEL:sec:simulation
For the scenarios considered in Section LABEL:sec:simulation, we present their true optimal values in Table A.1.
Table A.1: True optimal values in S1-S4 when $K=2$.
S1 S2 S3 S4
$\mathcal{V}^{*}_{1}$ $\mathcal{V}^{*}_{2}$ $\mathcal{V}^{*}_{1}$ $\mathcal{V}^{*}_{2}$ $\mathcal{V}^{*}_{1}$ $\mathcal{V}^{*}_{2}$ $\mathcal{V}^{*}_{1}$ $\mathcal{V}^{*}_{2}$
1.89 2.47 1.89 2.33 2.12 2.82 2.12 2.83
In addition to the results in Section LABEL:sec:simulation when $n=200$, we also present the case when $n=100$ in Figures A.1 and  A.2. The general conclusions are similar as before, but as the sample size decreases, the disagreement rate, RMSE and misclassification rate increases. Besides, FITR-Ramp has smaller RMSE compared to FITR-IntL in S3 and S4, which is different from the case when $n=200$. This may be due to the reason that FITR-IntL is unstable when the sample size is small and the variance is large.
We also explore the case where we have $K=3$ outcomes. Consider the following four scenarios. In all scenarios, the main effects are
$m_{1}(\bm{X})$ $=1+2X_{1}+X_{2}^{2}+X_{1}X_{2},$
$m_{2}(\bm{X})$ $=1+2X_{1}^{2}+1.5X_{2}+0.5X_{1}X_{2},$
$m_{3}(\bm{X})$ $=1+X_{1}+X_{2}.$
The noise terms are the same as that in Section A.2. The interaction terms are defined as follows:
  • In linear scenarios S5 and S6,
    $T_{1}(\bm{X},A)$ $=0.5A(0.2-X_{1}-2X_{2}),$
    $T_{2}(\bm{X},A)$ $=0.8A(0.2-X_{1}-\gamma_{1}X_{2}),$
    $T_{3}(\bm{X},A)$ $=0.6A(0.2-X_{1}-2.2X_{2}),$
    where $\gamma_{1}=1.8$ in S5 and $\gamma_{1}=1.4$ in S6.
  • In nonlinear scenarios S7 and S8,
    $T_{1}(\bm{X},A)$ $=1.0A(-2.2-e^{X_{1}}-e^{X_{2}}),$
    $T_{2}(\bm{X},A)$ $=1.5A(-\gamma_{2}-e^{X_{1}}-e^{X_{2}}),$
    $T_{3}(\bm{X},A)$ $=1.5A(-2.1-e^{X_{1}}-e^{X_{2}}),$
    where $\gamma_{2}=2.3$ in S7 and $\gamma_{2}=2.4$ in S8.
The first two outcomes are the same as in Section LABEL:sec:simulation when $K=2$. The true optimal values of each scenario is summarized in Table A.2.
Table A.2: True optimal values of S3 and S4 when $K=3$.
S1 S2 S3 S4
$\mathcal{V}^{*}_{1}$ $\mathcal{V}^{*}_{2}$ $\mathcal{V}^{*}_{3}$ $\mathcal{V}^{*}_{1}$ $\mathcal{V}^{*}_{2}$ $\mathcal{V}^{*}_{3}$ $\mathcal{V}^{*}_{1}$ $\mathcal{V}^{*}_{2}$ $\mathcal{V}^{*}_{3}$ $\mathcal{V}^{*}_{1}$ $\mathcal{V}^{*}_{2}$ $\mathcal{V}^{*}_{3}$
1.89 2.47 1.72 1.89 2.33 1.72 2.12 2.82 2.18 2.12 2.83 2.18
Figures A.3 and  A.4 contains the simulation results when $K=3$ and $n=200$, and Figures A.5 and  A.6 contains the simulation results when $K=3$ and $n=100$. When the number of outcomes $K$ increases, the disagreement rate, the RMSE and misclassification rate all decreases. Although the discrepancy between $R_{1}$ and $R_{2}$ is larger in S6 than in S5, the RMSE and misclassification rate of $R_{1}$ are almost the same as in S5. Remember that RMSE and misclassification rate are larger in S2 than in S1, although $R_{1}$ and $R_{2}$ are exactly the same as in S5 and S6 correspondingly. This reflects the advantage of using the third outcome, in spite of the fact that $R_{2}$ and $R_{3}$ differ from $R_{1}$ in opposite directions.
Figure A.1: The mean and standard deviation of disagreement rate between FITR $\widehat{f}_{1n}$ learned using SepL, FITR-Ramp or FITR-IntL, and the true secondary outcome ITR $f_{2}^{*}$ when $K=2$ and $n=100$. The dashed line represents the true disagreement rate $\mathbb{P}(f_{1}^{*}f_{2}^{*}<0)$.
(a) RMSE
(b) Accuracy
Figure A.2: (a) The RMSE of value functions and (b) the mean and standard deviation of accuracy when FITR is learned using SepL, FITR-Ramp or FITR-IntL when $K=2$ and $n=100$.
Figure A.3: The mean and standard deviation of disagreement rate between FITR $\widehat{f}_{1n}$ learned using SepL, FITR-Ramp or FITR-IntL, and the true secondary outcome ITRs $f_{2}^{*},f_{3}^{*}$ when $K=3$ and $n=200$. The subtitle of each subfigure refers to the secondary outcome for which we are estimating the disagreement rate. For example, “Outcome3” refers to the disagreement rate $\mathbb{P}(\widehat{f}_{1n}f_{3}^{*}<0)$. The dashed line represents the true disagreement rate $\mathbb{P}(f_{1}^{*}f_{2}^{*}<0)$ or $\mathbb{P}(f_{1}^{*}f_{3}^{*}<0)$.
(a) RMSE
(b) Accuracy
Figure A.4: (a) The RMSE of value functions and (b) the mean and standard deviation of accuracy when FITR is learned using SepL, FITR-Ramp or FITR-IntL when $K=3$ and $n=200$.
Figure A.5: The mean and standard deviation of disagreement rate between FITR $\widehat{f}_{1n}$ learned using SepL, FITR-Ramp or FITR-IntL, and the true secondary outcome ITRs $f_{2}^{*},f_{3}^{*}$ when $K=3$ and $n=100$. The subtitle of each subfigure refers to the secondary outcome for which we are estimating the disagreement rate. For example, “Outcome3” refers to the disagreement rate $\mathbb{P}(\widehat{f}_{1n}f_{3}^{*}<0)$. The dashed line represents the true disagreement rate $\mathbb{P}(f_{1}^{*}f_{2}^{*}<0)$ or $\mathbb{P}(f_{1}^{*}f_{3}^{*}<0)$.
(a) RMSE
(b) Accuracy
Figure A.6: (a) The RMSE of value functions and (b) the mean and standard deviation of accuracy when FITR is learned using SepL, FITR-Ramp or FITR-IntL when $K=3$ and $n=100$.
A.4 Additional Results for Sensitivity Analysis in Section LABEL:sec:simulation
This section contains the details of the sensitivity analysis in Section LABEL:sec:simulation. In Table A.3, we show different values of $\rho$ and their corresponding agreement rates between $f_{1}^{*}$ and $f_{2}^{*}$ when $n=200$ and $N_{2}=4n$.
Table A.3: The change of RMSE and accuracy when the similarity between outcomes is changed.
$\rho$ $\mathbb{P}(f_{1}^{*}f_{2}^{*}>0)$ Model $\mathbb{P}(\widehat{f}_{1n}f_{2}^{*}>0)$ RMSE $\frac{\text{RMSE}}{\text{RMSE}_{SepL}}$ Misclassification Rate
1 100% SepL 0.905 (0.026) 0.034 1.000 0.095 (0.026)
FITR-IntL 0.925 (0.024) 0.023 0.682 0.075 (0.024)
FITR-Ramp 0.936 (0.022) 0.019 0.569 0.064 (0.022)
0.9 98.57% SepL 0.904 (0.025) 0.034 1.000 0.095 (0.026)
FITR-IntL 0.922 (0.022) 0.023 0.679 0.076 (0.024)
FITR-Ramp 0.935 (0.022) 0.019 0.557 0.064 (0.022)
0.8 96.81% SepL 0.900 (0.025) 0.034 1.000 0.095 (0.026)
FITR-IntL 0.916 (0.023) 0.024 0.727 0.078 (0.025)
FITR-Ramp 0.928 (0.023) 0.021 0.616 0.070 (0.023)
0.7 94.48% SepL 0.891 (0.025) 0.034 1.000 0.095 (0.026)
FITR-IntL 0.909 (0.024) 0.026 0.770 0.081 (0.026)
FITR-Ramp 0.920 (0.026) 0.024 0.698 0.076 (0.023)
0.6 91.42% SepL 0.874 (0.027) 0.034 1.000 0.095 (0.026)
FITR-IntL 0.895 (0.027) 0.028 0.834 0.086 (0.026)
FITR-Ramp 0.908 (0.030) 0.027 0.811 0.085 (0.023)
0.5 87.56% SepL 0.847 (0.029) 0.034 1.000 0.095 (0.026)
FITR-IntL 0.873 (0.032) 0.033 0.970 0.094 (0.028)
FITR-Ramp 0.888 (0.038) 0.033 0.976 0.096 (0.025)
0.4 83.16% SepL 0.810 (0.031) 0.034 1.000 0.095 (0.026)
FITR-IntL 0.843 (0.040) 0.037 1.086 0.101 (0.028)
FITR-Ramp 0.858 (0.046) 0.039 1.169 0.106 (0.029)
A.5 Additional Results for Real Data Analysis in Section LABEL:sec:realdata
Two covariates were shown to be informative for tailoring treatments in a prior study (Chen et al.,, 2021). The first is the NEO-Five Factor Inventory score, where the NEO Personality Inventory is a 240-item measurement designed to assess personality in the domains of neuroticism, extraversion, openness, and so on; we focus on the neuroticism domain. The second informative measure is the Flanker Interference Accuracy score, where a higher value indicates reduced cognitive control. Five additional baseline variables are used, including sex, age, education years, Edinburgh Handedness Inventory (EHI) score, and the QIDS score at the beginning of the study.
We list the coefficients of the ITR $\widehat{f}_{\text{QIDS}}$ of the primary outcome estimated using SepL, FITR-IntL and FITR-Ramp in Table A.4, along with the secondary outcome ITRs $\widetilde{f}_{\text{CGI}}$ and $\widetilde{f}_{\text{SAS}}$ estimated using SepL. They are learned with the complete dataset and the linear kernel. We can conclude that the coefficients fitted by different methods are generally close, which suggests that the fusion penalty will not dramatically change an ITR compared to SepL. The coefficients are more similar for QIDS-change and CGI, which is expected since SAS measures the impacts of depression on social functioning. The Flanker variable has a different sign in the ITR of SAS than the other two variables.
Table A.4: Coefficients of the estimated ITRs of the three outcomes when linear kernel is used.
Outcome Model intercept sex age education EHI QIDS NEO Flanker
QIDS- change SepL 0.130 0.105 0.231 -0.126 0.066 0.031 0.251 -0.345
FITR-IntL 0.137 0.098 0.126 -0.020 0.032 0.065 0.087 -0.174
FITR-Ramp 0.063 0.038 0.078 -0.036 0.025 0.027 0.071 -0.108
CGI SepL 0.358 0.019 0.275 -0.043 0.086 0.014 0.155 -0.281
SAS SepL 0.207 0.014 0.147 -0.102 -0.051 0.202 0.157 0.029
Appendix B ADDITIONAL ASSUMPTIONS, RESULTS AND PROOFS FOR SECTION LABEL:sec:theory
In this section, we provide additional assumptions, theoretical results and proofs. The proofs are extended from outcome weighted learning with a hinge loss to our problem with logistic loss and nonconvex ramp loss. Besides, we also used a new change of measure technique in Section B.5, and derived a new result in Corollary B.2 under a relatively weak Assumption B.5.
B.1 Additional Assumptions and Theoretical Results
We present the details about the ignorability, consistency, and positivity assumptions, which are common in causal inference to identify the treatment effect.
Assumption B.1 (Ignorability).
The treatment $A$ is independent of the potential outcomes $R_{1}^{*}(a)$ given covariates $\bm{X}$.
Assumption B.2 (Consistency).
The observed outcome $R_{1}$ under a treatment $A=a$ equals the potential outcome $R_{1}(a)$ for all $a\in\mathcal{A}$.
Assumption B.3 (Positivity).
There exists $p_{0}>0$ such that $\pi(a;\bm{x})\equiv P(A=a|\bm{X}=\bm{x})>p_{0}$ for all $a\in\mathcal{A}$ and all $\bm{x}\in\mathcal{X}$.
To bound the risk with 0-1 loss using the risk with convex surrogate loss, we need another common assumption in classification literature, which assumes the noise conditions around the decision boundary (Bartlett et al.,, 2006). For Assumption B.4, it has been shown that the boundary assumption of $\eta(\bm{x})$ regarding $\beta$ is equivalent to the misclassification assumption of $f_{1}$ regarding $\alpha$ (Bartlett et al.,, 2006).
Assumption B.4.
Assume that the distribution of $\bm{X}$ satisfies the following Tsybakov’s noise assumption: there exists a constant $C>0$ such that for all sufficiently small $t>0$ we have
($\mathbb{P}(\left\{\bm{X}\in\mathcal{X}:\left\lvert 2\eta(\bm{X})-1\right\rvert\leq t\right\})\leq Ct^{\beta}$
for some $\beta>0$. Let $\alpha=\beta/(1+\beta)$ so that $\alpha\in(0,1]$.
Assumption B.5 is necessary when bounding the decision accuracy. It is a weak assumption that only requires one of the treatments to yield a positive mean reward, which is easily achievable by shifting the original rewards with a positive constant
Assumption B.5.
Suppose the conditional expectation of rewards satisfies $\sum_{a\in\mathcal{A}}r_{1}^{(a)}(\bm{x})\geq c_{r}$ for some constant $c_{r}>0$ for all $\bm{x}\in\mathcal{X}$.
Now we can present the convergence rates of the value function $\mathcal{V}_{1}(\widehat{f}_{1n})$ and the misclassification rate $\mathbb{P}(\widehat{f}_{1n}f_{1}^{*}<0)$ when $K=2$.
Theorem B.1.
Under Assumptions B.1-B.4 and LABEL:asp:bounded.reward-LABEL:asp:conv.secondary, when $K=2$, the value function of the estimated FITR satisfies
$\mathcal{V}_{1}(f_{1}^{*})-\mathcal{V}_{1}(\widehat{f}_{1n})\lesssim(\delta_{1n}(\tau)+\mu_{1n})^{\frac{1}{2-\alpha}},$ (B.1)
with probability at least $1-2e^{\tau}$.
Corollary B.2.
Under Assumptions B.1-B.5 and LABEL:asp:bounded.reward-LABEL:asp:conv.secondary, the misclassification rate satisfies
$\mathbb{P}(\widehat{f}_{1n}f_{1}^{*}<0)\lesssim(\delta_{1n}(\tau)+\mu_{1n})^{\frac{\alpha}{2-\alpha}}$ (B.2)
with probability at least $1-2e^{\tau}$.
Notice that the convergence rate $\widetilde{\delta}_{2N_{2}}(\tau)$ of $\widetilde{f}_{2}$ does not appear in (B.1), since the term containing $\widetilde{\delta}_{2N_{2}}(\tau)$ is not dominant in the proof due to the assumption that $\widetilde{\delta}_{2N_{2}}(\tau)=o(1)$. By comparing $\delta_{1n}(\tau)+\mu_{1n}$ and $\delta_{1n}^{(0)}(\tau)$, it can be concluded that FITR has a slower convergence rate for the value function than SepL. This is due to the fact that the fusion penalty introduces bias for maximizing the primary outcome.
The results in Theorem B.1 and Corollary B.2 can also be generalized to $K\geq 2$.
Corollary B.3.
Under Assumptions B.1-B.4 and LABEL:asp:bounded.reward-LABEL:asp:conv.secondary, when $K\geq 2$, the value function of the estimated FITR $\widehat{f}_{1n}$ satisfies
$\mathcal{V}_{1}(f_{1}^{*})-\mathcal{V}_{1}(\widehat{f}_{1n})\lesssim(\delta_{1n}(\tau)+\mu_{1n})^{\frac{1}{2-\alpha}},$ (B.3)
with probability at least $1-Ke^{\tau}$.
Under Assumptions B.1-B.5 and LABEL:asp:bounded.reward-LABEL:asp:conv.secondary, the misclassification rate satisfies
$\mathbb{P}(\widehat{f}_{1n}f_{1}^{*}<0)\lesssim(\delta_{1n}(\tau)+\mu_{1n})^{\frac{\alpha}{2-\alpha}}$ (B.4)
with probability at least $1-Ke^{\tau}$.
Remark.
If $\widetilde{f}_{k}$ is the sign of the decision function learned by SepL with sample size $N_{k}$, we can directly use Corollary B.2 with $\mu_{kN_{k}}=0$ to find that $\widetilde{\delta}_{kN_{k}}(\tau)$ in Assumption LABEL:asp:conv.secondary can be expressed as $\widetilde{\delta}_{kN_{k}}(\tau)=(\delta_{kN_{k}}^{(0)}(\tau))^{\frac{\alpha}{2-\alpha}}$, where
$\delta_{kN_{k}}^{(0)}(\tau):=\lambda_{kN_{k}}^{-\frac{1}{2}}N_{k}^{-\frac{1}{2}}\left[\sqrt{\tau}+\sigma_{kN_{k}}^{(1-\nu/2)(1+\delta)d}\right]+\lambda_{kN_{k}}\sigma_{kN_{k}}^{d}+(2d)^{qd/2}\sigma_{kN_{k}}^{-qd}.$
Remark.
With a stronger assumption $\lvert r_{1}^{(1)}(\bm{x})-r_{1}^{(-1)}(\bm{x})\rvert>0$ for all $\bm{x}$, which assumes that the treatment effect is nonzero for all patients, $\mathbb{P}(\widehat{f}_{1n}f_{1}^{*}<0)$ can be directly bounded by a multiple of $\mathcal{V}_{1}(f_{1}^{*})-\mathcal{V}_{1}(\widehat{f}_{1n})$ in Theorem B.1. Corollary B.2 does not rely on the assumption and only assumes that $r_{1}^{(a)}(\bm{x})$ cannot be zero at the same time for all $a\in\mathcal{A}$. Under the weaker assumption, the convergence rate of $\mathbb{P}(\widehat{f}_{1n}f_{1}^{*}<0)$ is slower with a scaling constant $\alpha\in(0,1]$.
B.2 Additional Details for Section LABEL:sec:theory
The parameters $\lambda_{1n},\mu_{1n},\kappa_{1n}$ should balance the estimation error and approximation error to achieve the minimal upper bound $\delta_{1n}(\tau)$. The minimal approximation error is $\left(\mu_{1n}\kappa_{1n}^{-1}\right)^{\frac{1}{1+q}}\lambda_{1n}^{\frac{q}{1+q}}$ when $\sigma_{1n}=\left(\mu_{1n}\kappa_{1n}^{-1}\lambda_{1n}^{-1}\right)^{\frac{1}{(1+q)d}}$. Then the best $\lambda_{1n}$ that balances the estimation error and the approximation error is $\lambda_{1n}=\left(\mu_{1n}\kappa_{1n}^{-1}\right)^{\frac{2(\Delta-1)+2(1+q)\nu}{3q+2\Delta+1}}n^{-\frac{1+q}{3q+2\Delta+1}}$. When $\mu_{1n}=0$, the convergence rate of the risk is
$\begin{split}\delta_{1n}^{(0)}(\tau):=\lambda_{1n}^{-\frac{1}{2}}n^{-\frac{1}{2}}\left[\sqrt{\tau}+\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\right]+\lambda_{1n}\sigma_{1n}^{d}+(2d)^{qd/2}\sigma_{1n}^{-qd}.\end{split}$
The minimum approximation error $\lambda_{1n}^{\frac{q}{1+q}}$ is obtained when $\sigma_{1n}=\lambda_{1n}^{-\frac{1}{(1+q)d}}$. Then the best $\lambda_{1n}$ is $n^{-\frac{1+q}{3q+2\Delta+1}}$.
To bound the agreement rate without the fusion penalty, first note that
$\mathbb{P}(\widehat{f}_{1n}f_{2}^{*}<0)=\mathbb{P}(\widehat{f}_{1n}f_{1}^{*}<0,f_{1}^{*}f_{2}^{*}>0)+\mathbb{P}(\widehat{f}_{1n}f_{1}^{*}>0,f_{1}^{*}f_{2}^{*}<0)\leq\mathbb{P}(\widehat{f}_{1n}f_{1}^{*}<0)+\mathbb{P}(f_{1}^{*}f_{2}^{*}<0),$
which allows us to obtain the bound with the help of the misclassification rate. For SepL with $\mu_{1n}=0$, we have $\mathbb{P}(\widehat{f}_{1n}f_{1}^{*}<0)\lesssim n^{-\frac{\alpha}{2-\alpha}\frac{q}{3q+2\Delta+1}}$ with probability at least $1-2e^{\tau}$ according to (LABEL:equ:min.delta.n0) and Corollary B.2. Therefore, we can conclude that for SepL
$\begin{split}\mathbb{P}(f_{1}^{*}f_{2}^{*}>0)-\mathbb{P}(\widehat{f}_{1n}f_{2}^{*}>0)\leq\mathbb{P}(\widehat{f}_{1n}f_{1}^{*}<0)\lesssim n^{-\frac{\alpha}{2-\alpha}\frac{q}{3q+2\Delta+1}}.\end{split}$ (B.5)
B.3 Proof of Lemma LABEL:lem:surrogate.risk
Proof.
First note that
$\mathcal{R}(\widehat{f}_{1n})-\mathcal{R}(f_{1}^{*})\leq\mathcal{R}(\widehat{f}_{1n})-\mathcal{R}(f_{1}^{*})+\lambda_{1n}\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}$ (B.6)
$\leq$ $\left\{\left[\mathcal{R}(\widehat{f}_{1n})+\lambda_{1n}\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}\right]-\inf_{f_{1}\in\mathcal{H}}\left[\mathcal{R}(f_{1})+\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}\right]\right\}$ (B.7)
+$+\left\{\inf_{f_{1}\in\mathcal{H}}\left[\mathcal{R}(f_{1})+\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}\right]-\mathcal{R}(f_{1}^{*})\right\}.$ (B.8)
Define $f_{1}^{\dagger}$ as the minimizer of $\mathcal{R}(f_{1})+\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}$ in $\mathcal{H}$. We will bound the two terms on the right-hand side separately.
To bound (B.8), we follow the construction in the proof of Theorem of 2.7 in Steinwart and Scovel, (2007). When $\mathcal{X}$ is the closed unit ball, on $\acute{\mathcal{X}}:=3\mathcal{X}$ define
$\acute{\eta}(\bm{x})=\begin{cases}\eta(\bm{x}),&\text{if }\left\lVert\bm{x}\right\rVert_{2}\leq 1,\\ \eta(\bm{x}/\left\lVert\bm{x}\right\rVert_{2}),&\text{otherwise}.\end{cases}$
Besides, let $\acute{\mathcal{X}}_{-1}:=\{x\in\acute{\mathcal{X}}:\acute{\eta}(\bm{x})<\frac{1}{2}\}$ and $\acute{\mathcal{X}}_{1}:=\{x\in\acute{\mathcal{X}}:\acute{\eta}(\bm{x})>\frac{1}{2}\}$. Fix a measurable $\acute{f}_{1}:\acute{\mathcal{X}}\mapsto[-1,1]$ that satisfies $\acute{f}_{1}=1$ on $\acute{\mathcal{X}}_{1}$, $\acute{f}_{1}=-1$ on $\acute{\mathcal{X}}_{-1}$ and $\acute{f}_{1}=0$ otherwise. The linear operator $V_{\sigma_{1n}}:L_{2}(\mathbb{R}^{d})\mapsto\mathcal{H}_{\sigma_{1n}}(\mathbb{R}^{d})$ defined by
[…]
is an isometric isomorphism (Steinwart et al.,, 2006). Consequently, we have
$\begin{split}&\inf_{f_{1}\in\mathcal{H}}\left[\mathcal{R}(f_{1})+\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}\right]-\mathcal{R}(f_{1}^{*})\\ \leq&\inf_{g\in L_{2}(\mathbb{R}^{d})}\left[\mathbb{E}(\ell_{1}\circ V_{\sigma_{1n}}g-\ell_{1}\circ f_{1}^{*})+\mathbb{E}(\ell_{2}\circ V_{\sigma_{1n}}g-\ell_{2}\circ f_{1}^{*})+\lambda_{1n}\lVert g\rVert_{L_{2}(\mathbb{R}^{d})}^{2}\right].\end{split}$
Now take a specific $g:=\big{(}\frac{\sigma_{1n}^{2}}{\pi}\big{)}^{d/4}\acute{f}_{1}$, and we obtain
$\lVert g\rVert_{L_{2}(\mathbb{R}^{d})}\leq\left(\frac{81\sigma_{1n}^{2}}{\pi}\right)^{d/4}\theta(d),$ (B.9)
where $\theta(d)$ denotes the volume of $\mathcal{X}$. Since $\-1\leq\acute{f}_{1}\leq 1$, it can be easily seen that $\-1\leq V_{\sigma_{1n}}g\leq 1$. Note that $\left\lvert\phi^{\prime}(t)\right\rvert=\left\lvert-\frac{e^{-t}}{1+e^{-t}}\right\rvert\leq 1$, so $\ell_{1}$ is Lipschitz continuous with respect to $f_{1}$ with Lipschitz constant $r/p_{0}$. It has been shown in Steinwart and Scovel, (2007) that
$\left\lvert V_{\sigma_{1n}}g(\bm{x})-f_{1}^{*}(\bm{x})\right\rvert\leq 8e^{-\sigma_{1n}^{2}\omega_{\bm{x}}^{2}/2d}.$
Therefore, Assumption LABEL:asp:geometric.noise for $t=2d/\sigma_{1n}^{2}$ yield
$\mathbb{E}(\ell_{1}\circ V_{\sigma_{1n}}g-\ell_{1}\circ f_{1}^{*})\lesssim\mathbb{E}\left\lvert V_{\sigma_{1n}}g-f_{1}^{*}\right\rvert\lesssim\mathbb{E}e^{-\sigma_{1n}^{2}\omega_{\bm{x}}^{2}/2d}\lesssim(2d)^{qd/2}\sigma_{1n}^{-qd}.$ (B.10)
Since $\ell_{2}$ is Lipschitz continuous with respect to $f_{1}$ with Lipschitz constant $\frac{\mu_{1n}\Omega_{12}}{\kappa_{1n}}\lVert\widetilde{f}_{2}\rVert_{\infty}$,
$\mathbb{E}(\ell_{2}\circ V_{\sigma_{1n}}g-\ell_{2}\circ f_{1}^{*})\lesssim\frac{\mu_{1n}}{\kappa_{1n}}\mathbb{E}\left\lvert V_{\sigma_{1n}}g-f_{1}^{*}\right\rvert\lesssim\frac{\mu_{1n}}{\kappa_{1n}}(2d)^{qd/2}\sigma_{1n}^{-qd}$ (B.11)
when $\lVert\widetilde{f}_{2}\rVert_{\infty}=1$. Combining (B.9), (B.10) and (B.11), we can bound the approximation error (B.8) as
$\inf_{f_{1}\in\mathcal{H}}\left[\mathcal{R}(f_{1})+\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}\right]-\mathcal{R}(f_{1}^{*})\lesssim\lambda_{1n}\sigma_{1n}^{d}+\left(1+\frac{\mu_{1n}}{\kappa_{1n}}\right)(2d)^{qd/2}\sigma_{1n}^{-qd}.$ (B.12)
To bound (B.7), we will use Talagrand’s inequality quoted as follows (Steinwart and Scovel,, 2007, Theorem 5.3).
Theorem B.4.
Let $\mathcal{H}$ be a set of bounded measurable functions from $\bm{Z}$ to $\mathbb{R}$, which is separable with respect to $\left\lVert\cdot\right\rVert_{\infty}$ and satisfies $\mathbb{E}h=0$ for all $h\in\mathcal{H}$. Furthermore, let $B>0$ and $b\geq 0$ be constants with $\left\lVert h\right\rVert_{\infty}\leq B$ and $\mathbb{E}h^{2}\leq b$ for all $h\in\mathcal{H}$. Then for all $\tau\geq 1$ and all $n\geq 1$ we have
$\mathbb{P}\left(\sup_{h\in\mathcal{H}}\mathbb{P}_{n}h>3\mathbb{E}\sup_{h\in\mathcal{H}}\mathbb{P}_{n}h+\sqrt{\frac{2\tau b}{n}}+\frac{B\tau}{n}\right)\leq e^{-\tau}.$
We first obtain a bound for $\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}$. Since $\mathbb{P}_{n}(\ell_{1}\circ\widehat{f}_{1n}+\ell_{2}\circ\widehat{f}_{1n})+\lambda_{1n}\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}\leq\mathbb{P}_{n}(\ell_{1}\circ f_{1}+\ell_{2}\circ f_{1})+\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}$ for any $f_{1}\in\mathcal{H}$, when taking $f_{1}=0$ we have
$\lambda_{1n}\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}\leq\mathbb{P}_{n}(\ell_{1}\circ f_{1}+\ell_{2}\circ f_{1})\leq\frac{r}{p_{0}}+\mu_{1n}\Omega_{12}.$
Since $r/p_{0}+\mu_{1n}\Omega_{12}\simeq M$, $\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}$ is bounded by $\sqrt{M/\lambda_{1n}}$, where $M:=r/p_{0}$. To this end, it suffices to consider the ball of radius $\sqrt{M/\lambda_{1n}}$. Therefore, the function class that we consider here is
$\mathcal{G}:=\left\{\ell_{1}\circ f_{1}+\ell_{2}\circ f_{1}+\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}-\left[\ell_{1}\circ f_{1}^{\dagger}+\ell_{2}\circ f_{1}^{\dagger}+\lambda_{1n}\lVert f_{1}^{\dagger}\rVert_{\mathcal{H}}^{2}\right]:f\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})\right\},$
where $B_{\mathcal{H}}(r)$ is the ball in $\mathcal{H}$ of radius $r$. Since $\ell_{1}$ is Lipschitz continuous with respect to $f_{1}$ and $\left\lVert f\right\rVert_{\infty}\leq\left\lVert f\right\rVert_{\mathcal{H}}$ for any $g\in\mathcal{G}$,
$\begin{split}\left\lvert g\right\rvert\leq&\left\lvert\ell_{1}\circ f_{1}-\ell_{1}\circ f_{1}^{\dagger}\right\rvert+\left\lvert\ell_{2}\circ f_{1}-\ell_{2}\circ f_{1}^{\dagger}\right\rvert+\lambda_{1n}\left\lvert\lVert f_{1}\rVert_{\mathcal{H}}^{2}-\lVert f_{1}^{\dagger}\rVert_{\mathcal{H}}^{2}\right\rvert\\ \leq&M\left\lvert f_{1}-f_{1}^{\dagger}\right\rvert+\mu_{1n}\Omega_{12}+M\\ \leq&2M\sqrt{M/\lambda_{1n}}+\mu_{1n}\Omega_{12}+M.\end{split}$
Hence, with $B:=2M\sqrt{M/\lambda_{1n}}+\mu_{1n}\Omega_{12}+M\simeq\lambda_{1n}^{-1/2}$, we have $\left\lVert g\right\rVert_{\infty}\leq B$.
Define the modulus of continuity of $\mathcal{G}$ by
$\omega_{n}(\mathcal{G},\epsilon):=\mathbb{E}\left(\sup_{g\in\mathcal{G},\mathbb{E}g^{2}\leq\epsilon}\left\lvert\mathbb{E}g-\mathbb{P}_{n}g\right\rvert\right),\quad\epsilon>0,$
where the supremum is measurable by the separability assumption on $\mathcal{G}$. Define the function class
:=$\mathcal{E}:=\left\{\mathbb{E}g-g:g\in\mathcal{G}\right\},$ (B.13)
and then we have $\omega_{n}(\mathcal{G},4B^{2})\geq\mathbb{E}\sup_{h\in\mathcal{E}}\mathbb{P}_{n}h$ since $\left\lvert\mathbb{E}g-g\right\rvert\leq 2B$. By Theorem B.4 we obtain
$\mathbb{P}\left(\sup_{h\in\mathcal{E}}\mathbb{P}_{n}h>3\omega_{n}(\mathcal{G},4B^{2})+\sqrt{\frac{2\tau(4B)^{2}}{n}}+\frac{2B\tau}{n}\right)\leq e^{-\tau}.$ (B.14)
Let $\bm{\epsilon}=\{\epsilon_{i}\}_{i=1}^{n}$ be a sequence of i.i.d. Rademacher variable. Then the local Rademacher average of $\mathcal{G}$ is defined by
$\operatorname{Rad}(\mathcal{G},n,\epsilon):=\mathbb{E}_{\bm{z}}\mathbb{E}_{\bm{\epsilon}}\sup_{g\in\mathcal{G},\mathbb{E}g^{2}\leq\epsilon}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}g(\bm{z}_{i})\right\rvert.$
It has been shown that
$\omega_{n}(\mathcal{G},\epsilon)\leq 2\operatorname{Rad}(\mathcal{G},n,\epsilon),\quad\epsilon>0$
by symmetrization (van der Vaart and Wellner,, 1996). Since
$\operatorname{Rad}(\mathcal{G},n,\epsilon)=B\operatorname{Rad}(B^{-1}\mathcal{G},n,B^{-2}\epsilon)$
for any $a>0$ by equation (37) of Steinwart and Scovel, (2007), we only need to obtain a bound for $\operatorname{Rad}(B^{-1}\mathcal{G},n,B^{-2}\epsilon)$. To this end, we will use Proposition 5.5 of Steinwart and Scovel, (2007) to bound the local Rademacher average, quoted as follows.
Theorem B.5.
Let $\mathcal{F}$ be a class of measurable functions from $\bm{Z}$ to $[-1,1]$ which is separable with respect to $\left\lVert\cdot\right\rVert_{\infty}$. Assume there are constants $a>0$ and $20<p<$ 2 with
$\sup_{\mathbb{P}_{n}}\log N(\epsilon,\mathcal{F},L_{2}(\mathbb{P}_{n}))\leq a\epsilon^{-p}$
for all $\epsilon>0$. Then there exists a constant $c_{p}>0$ depending only on $p$ such that for all $n\geq 1$ and all $\epsilon>0$ we have
$\operatorname{Rad}(\mathcal{F},n,\epsilon)\leq c_{p}\max\left\{\epsilon^{\frac{1}{2}-\frac{p}{4}}\left(\frac{a}{n}\right)^{\frac{1}{2}},\left(\frac{a}{n}\right)^{\frac{2}{2+p}}\right\}.$
Now we need to find some constants $a>0$ and $20<p<$ 2 such that
$\sup_{\mathbb{P}_{n}}\log N(\epsilon,B^{-1}\mathcal{G},L_{2}(\mathbb{P}_{n}))\leq a\epsilon^{-p}$
for some $a\geq 1,0<p<2$ and for all $\epsilon>0$. To this end, note that
$\begin{split}&\log N(\epsilon,B^{-1}\mathcal{G},L_{2}(\mathbb{P}_{n}))\\ =&\log N\left(B^{-1}\left\{\ell_{1}\circ f_{1}+\ell_{2}\circ f_{1}+\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}:f\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})\right\},\epsilon,L_{2}(\mathbb{P}_{n})\right)\\ \leq&\log N\left(B^{-1}\left\{\ell_{1}\circ f_{1}+\ell_{2}\circ f_{1}:f\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})\right\},\epsilon,L_{2}(\mathbb{P}_{n})\right)\\ &+\log N\left(B^{-1}\left\{\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}:f\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})\right\},\epsilon,L_{2}(\mathbb{P}_{n})\right)\end{split}$
by the subadditivity of the entropy. For the first term on the right-hand side, for any $f_{1},f_{1}^{\prime}\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})$, let $u:=B^{-1}(\ell_{1}+\ell_{2})\circ f_{1}$ and $u^{\prime}:=B^{-1}(\ell_{1}+\ell_{2})\circ f_{1}^{\prime}$. Since $\ell_{1}$ and $\ell_{2}$ are Lipschitz continuous with respect to $f_{1}$,
$\left\lVert u-u^{\prime}\right\rVert_{L_{2}(\mathbb{P}_{n})}\leq B^{-1}\left(M+\frac{\mu_{1n}\Omega_{12}}{\kappa_{1n}}\right)\lVert f-f^{\prime}\rVert_{L_{2}(\mathbb{P}_{n})}.$
With $u,u^{\prime}\in B^{-1}\left\{\ell_{1}\circ f_{1}+\ell_{2}\circ f_{1}:f\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})\right\}$,
$\begin{split}&\log N\left(B^{-1}\left\{\ell_{1}\circ f_{1}+\ell_{2}\circ f_{1}:f\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})\right\},\epsilon,L_{2}(\mathbb{P}_{n})\right)\\ \leq&\log N\left(B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}}),\frac{B\epsilon}{M+\mu_{1n}\Omega_{12}\kappa_{1n}^{-1}},L_{2}(\mathbb{P}_{n})\right)\\ \leq&\log N\left(B_{\mathcal{H}},\gamma_{n}\epsilon,L_{2}(\mathbb{P}_{n})\right),\end{split}$
where
$\gamma_{n}:=\frac{B}{\sqrt{M/\lambda_{1n}}(M+\mu_{1n}\Omega_{12}\kappa_{1n}^{-1})}\simeq\frac{1}{1+\mu_{1n}\kappa_{1n}^{-1}}$
For the second term on the right-hand side, it follows that
$\log N\left(B^{-1}\left\{\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}:f\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})\right\},\epsilon,L_{2}(\mathbb{P}_{n})\right)\leq\log\frac{M}{B\epsilon}$
since $\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}\leq M$ for all $f\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})$. Therefore, we can conclude that
$\log N(\epsilon,B^{-1}\mathcal{G},L_{2}(\mathbb{P}_{n}))\leq\log N\left(B_{\mathcal{H}},\gamma_{n}\epsilon,L_{2}(\mathbb{P}_{n})\right)+\log\frac{M}{B\epsilon}.$
Theorem 2.1 of Steinwart and Scovel, (2007) then yields that
$\sup_{\mathbb{P}_{n}}\log N(\epsilon,B^{-1}\mathcal{G},L_{2}(\mathbb{P}_{n}))\lesssim\sigma_{1n}^{(1-\nu/2)(1+\delta)d}(\gamma_{n}\epsilon)^{-\nu},$
where $\sigma_{1n}>0$ is the parameter of the Gaussian kernel associated with $\mathcal{H}$, and $00<\nu\leq 2,\delta>0,\epsilon>$ 0. Therefore, we have $a=\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{n}^{-\nu}$, $p=\nu$ and
$\operatorname{Rad}(\mathcal{G},n,\epsilon)\leq c_{p}\max\left\{B^{\frac{p}{2}}\epsilon^{\frac{1}{2}-\frac{p}{4}}\left(\frac{a}{n}\right)^{\frac{1}{2}},B\left(\frac{a}{n}\right)^{\frac{2}{2+p}}\right\}.$
With $\epsilon=4B^{2}$, we can bound the modulus of continuity as
$\omega_{n}(\mathcal{G},\epsilon)\leq 2\operatorname{Rad}(\mathcal{G},n,\epsilon)\lesssim Ba^{\frac{1}{2}}n^{-\frac{1}{2}}\simeq\lambda_{1n}^{-\frac{1}{2}}\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{n}^{-\nu}n^{-\frac{1}{2}}.$ (B.15)
The definition of $\widehat{f}_{1n}$ yields that
$\mathbb{P}_{n}\left\{\ell_{1}\circ\widehat{f}_{1n}+\ell_{2}\circ\widehat{f}_{1n}+\lambda_{1n}\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}-\left[\ell_{1}\circ f_{1}^{\dagger}+\ell_{2}\circ f_{1}^{\dagger}+\lambda_{1n}\lVert f_{1}^{\dagger}\rVert_{\mathcal{H}}^{2}\right]\right\}\leq 0.$
Therefore,
$\begin{split}&\left[\mathcal{R}(\widehat{f}_{1n})+\lambda_{1n}\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}\right]-\left[\mathcal{R}(f_{1}^{\dagger})+\lambda_{1n}\lVert f_{1}^{\dagger}\rVert_{\mathcal{H}}^{2}\right]\\ =&\mathbb{E}\left\{\ell_{1}\circ\widehat{f}_{1n}+\ell_{2}\circ\widehat{f}_{1n}+\lambda_{1n}\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}-\left[\ell_{1}\circ f_{1}^{\dagger}+\ell_{2}\circ f_{1}^{\dagger}+\lambda_{1n}\lVert f_{1}^{\dagger}\rVert_{\mathcal{H}}^{2}\right]\right\}\\ \leq&(\mathbb{E}-\mathbb{P}_{n})\left\{\ell_{1}\circ\widehat{f}_{1n}+\ell_{2}\circ\widehat{f}_{1n}+\lambda_{1n}\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}-\left[\ell_{1}\circ f_{1}^{\dagger}+\ell_{2}\circ f_{1}^{\dagger}+\lambda_{1n}\lVert f_{1}^{\dagger}\rVert_{\mathcal{H}}^{2}\right]\right\}\\ \leq&\sup_{h\in\mathcal{E}}\mathbb{P}_{n}h,\end{split}$
where $\mathcal{E}$ is defined in (B.13). Note that $\left\lVert h\right\rVert_{\infty}\leq 2B\lesssim\lambda_{1n}^{-\frac{1}{2}}$ and $\mathbb{E}h^{2}\leq 4B^{2}\lesssim\lambda_{1n}^{-1}$ for all $h\in\mathcal{E}$. Plugging (B.15) into (B.14) and we have
$\begin{split}&\left[\mathcal{R}(\widehat{f}_{1n})+\lambda_{1n}\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}\right]-\left[\mathcal{R}(f_{1}^{\dagger})+\lambda_{1n}\lVert f_{1}^{\dagger}\rVert_{\mathcal{H}}^{2}\right]\leq\sup_{h\in\mathcal{E}}\mathbb{P}_{n}h\\ \lesssim&\lambda_{1n}^{-\frac{1}{2}}\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{n}^{-\nu}n^{-\frac{1}{2}}+\sqrt{\frac{2\tau\lambda_{1n}^{-1}}{n}}+\frac{\tau\lambda_{1n}^{-\frac{1}{2}}}{n}\\ \lesssim&\lambda_{1n}^{-\frac{1}{2}}n^{-\frac{1}{2}}\left[\sqrt{\tau}+\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{n}^{-\nu}\right]\end{split}$ (B.16)
with probability at least $1-e^{-\tau}$ for any $\tau\geq 1$.
Finally, plug the upper bounds (B.16) and (B.12) into (B.7) and (B.8) and we get the results. ∎
B.4 Proof of Theorem LABEL:thm:convergence.rate.disagree
Proof.
For $\ell_{2}$ with the ramp loss, note that $\psi_{\kappa_{n}}(f_{1}^{*}f_{2}^{*})={\mathds{1}}\{f_{1}^{*}f_{2}^{*}<0\}$ since $f_{1}^{*}f_{2}^{*}$ takes values only in $\{-1,1\}$ when $\kappa_{n}\leq 1$. Besides, $\psi_{\kappa_{n}}(f_{1}f_{2}^{*})\geq{\mathds{1}}\{f_{1}f_{2}^{*}<0\}$ for any $f_{1}$ by the definition of the ramp loss $\psi$. Hence, we obtain the relationship between the excess risks under the ramp loss and the 0-1 loss as
$\mathbb{E}\psi_{\kappa_{n}}(f_{1}f_{2}^{*})-\mathbb{E}\psi_{\kappa_{n}}(f_{1}^{*}f_{2}^{*})\geq\mathbb{E}{\mathds{1}}\{f_{1}f_{2}^{*}<0\}-\mathbb{E}{\mathds{1}}\{f_{1}^{*}f_{2}^{*}<0\}.$ (B.17)
Since $\widetilde{f}_{2}$ and $f_{2}^{*}$ are binary decision functions,
$\lvert\psi_{\kappa_{n}}(f_{1}\widetilde{f}_{2})-\psi_{\kappa_{n}}(f_{1}f_{2}^{*})\rvert={\mathds{1}}\{\widetilde{f}_{2}f_{2}^{*}<0\}\lvert\psi_{\kappa_{n}}(f_{1}\widetilde{f}_{2})-\psi_{\kappa_{n}}(f_{1}f_{2}^{*})\rvert\leq{\mathds{1}}\{\widetilde{f}_{2}f_{2}^{*}<0\}\leq\widetilde{\delta}_{2N_{2}}(\tau)$ (B.18)
with probability at least $1-e^{\tau}$ for any $f_{1}$ by Assumption LABEL:asp:conv.secondary. The first inequality comes from the fact that one of $\psi_{\kappa_{n}}(f_{1}\widetilde{f}_{2})$ and $\psi_{\kappa_{n}}(f_{1}f_{2}^{*})$ must be zero and the other is bounded by one. Therefore, we have
$\begin{split}\mathbb{E}\ell_{2}\circ f_{1}-\mathbb{E}\ell_{2}\circ f_{1}^{*}=&\mu_{1n}\Omega_{12}\left[\mathbb{E}\psi_{\kappa_{n}}(f_{1}\widetilde{f}_{2})-\mathbb{E}\psi_{\kappa_{n}}(f_{1}^{*}\widetilde{f}_{2})\right]\\ \geq&\mu_{1n}\Omega_{12}\left[\mathbb{E}\psi_{\kappa_{n}}(f_{1}f_{2}^{*})-\mathbb{E}\psi_{\kappa_{n}}(f_{1}^{*}f_{2}^{*})-2\widetilde{\delta}_{2N_{2}}(\tau)\right]\\ \geq&\mu_{1n}\Omega_{12}\left[\mathbb{E}{\mathds{1}}\{f_{1}f_{2}^{*}<0\}-\mathbb{E}{\mathds{1}}\{f_{1}^{*}f_{2}^{*}<0\}-2\widetilde{\delta}_{2N_{2}}(\tau)\right],\end{split}$ (B.19)
where the first inequality comes from (B.18) and the second inequality comes from (B.17).
Since $\mathbb{E}\ell_{1}\circ\widehat{f}_{1n}-\mathbb{E}\ell_{1}\circ f_{1}^{*}+\mathbb{E}\ell_{2}\circ\widehat{f}_{1n}-\mathbb{E}\ell_{2}\circ f_{1}^{*}\lesssim\delta_{1n}(\tau)$, we have
$\mu_{1n}[\mathbb{E}{\mathds{1}}\{\widehat{f}_{1n}f_{2}^{*}<0\}-\mathbb{E}{\mathds{1}}\{f_{1}^{*}f_{2}^{*}<0\}-\widetilde{\delta}_{2N_{2}}(\tau)]\lesssim\mathbb{E}\ell_{2}\circ\widehat{f}_{1n}-\mathbb{E}\ell_{2}\circ f_{1}^{*}\lesssim\delta_{1n}(\tau)$
with probability at least $1-2e^{\tau}$, and the results in Theorem LABEL:thm:convergence.rate.disagree follows. ∎
B.5 Proof of Theorem B.1
Proof.
Define $\widehat{U}_{1n}:=\mathbb{E}{\mathds{1}}\{\widehat{f}_{1n}f_{1}^{*}<0\}$ and $\Delta\mathcal{V}_{1}(f_{1}):=\mathcal{V}_{1}(f_{1}^{*})-\mathcal{V}_{1}(f_{1})$ for simplicity.
To utilize existing results in general classification problems, we can rewrite our loss functions with a change of measure. Let $h(\cdot)$ be the probability distribution function of the covariates $\bm{X}$. Then the expectation of $\ell_{1}$ can be written as
$\mathbb{E}\ell_{1}\circ f_{1}=\mathbb{E}\left[\frac{R_{1}}{\pi(A;\bm{X})}\phi(Af_{1}(\bm{X}))\right]=\int_{\mathcal{X}_{+}}\sum_{a\in\mathcal{A}}\frac{r_{1}^{(a)}(\bm{x})}{\pi(a;\bm{x})}\phi(af_{1}(\bm{x}))\pi(a;\bm{x})h(\bm{x})d\bm{x},$
[…]
$\pi^{\prime}(a;\bm{x}):=\begin{cases}\frac{r_{1}^{(a)}(\bm{x})}{g(\bm{x})},&\text{if }g(\bm{x})>0\\ \frac{1}{2},&\text{otherwise},\end{cases}$
so that $\pi^{\prime}\in[0,1]$ by Assumption LABEL:asp:bounded.reward can be regarded as a new policy for sampling the treatments. Then we obtain
$\begin{split}\mathbb{E}\ell_{1}\circ f_{1}=&C_{R_{1}}\int_{\mathcal{X}_{+}}\sum_{a\in\mathcal{A}}\phi(af_{1}(\bm{x}))\frac{r_{1}^{(a)}(\bm{x})}{g(\bm{x})}\frac{g(\bm{x})h(\bm{x})}{C_{R_{1}}}d\bm{x}\\ =&C_{R_{1}}\int\sum_{a\in\mathcal{A}}\phi(af_{1}(\bm{x}))\pi^{\prime}(a;\bm{x})h^{\prime}(\bm{x})d\bm{x}.\end{split}$
Denote $\mathbb{E}^{\prime}$ as the expectation corresponding to the distributions $h^{\prime}$ and $\pi^{\prime}$, so we get
$\mathbb{E}\ell_{1}\circ f_{1}=C_{R_{1}}\mathbb{E}^{\prime}\phi(Af_{1}).$
Conversely, for the 0-1 loss, the difference between value functions can be written as
$\begin{split}\Delta\mathcal{V}_{1}(f_{1})=&\mathcal{V}_{1}(f_{1}^{*})-\mathcal{V}_{1}(f_{1})\\ =&\mathbb{E}\left[\frac{R_{1}}{\pi(A;\bm{X})}{\mathds{1}}\{Af_{1}(\bm{X})<0\}\right]-\mathbb{E}\left[\frac{R_{1}}{\pi(A;\bm{X})}{\mathds{1}}\{Af_{1}^{*}(\bm{X})<0\}\right]\\ =&C_{R_{1}}\left[\mathbb{E}^{\prime}{\mathds{1}}\{Af_{1}<0\}-\mathbb{E}^{\prime}{\mathds{1}}\{Af_{1}^{*}<0\}\right].\end{split}$
By Theorem 3 of Bartlett et al., (2006) and Assumption B.4,
$\begin{split}&\mathbb{E}\ell_{1}\circ f_{1}-\mathbb{E}\ell_{1}\circ f_{1}^{*}=C_{R_{1}}\mathbb{E}^{\prime}\phi(Af_{1})-C_{R_{1}}\mathbb{E}^{\prime}\phi(Af_{1}^{*})\\ \geq&C_{R_{1}}c\left[\mathbb{E}^{\prime}{\mathds{1}}\{Af_{1}<0\}-\mathbb{E}^{\prime}{\mathds{1}}\{Af_{1}^{*}<0\}\right]^{\alpha}\rho\left(\frac{\left[\mathbb{E}^{\prime}{\mathds{1}}\{Af_{1}<0\}-\mathbb{E}^{\prime}{\mathds{1}}\{Af_{1}^{*}<0\}\right]^{1-\alpha}}{2c}\right)\\ =&C_{R_{1}}c\left[\frac{1}{C_{R_{1}}}\Delta\mathcal{V}_{1}(f_{1})\right]^{\alpha}\rho\left[\frac{1}{2c}\left(\frac{1}{C_{R_{1}}}\Delta\mathcal{V}_{1}(f_{1})\right)^{1-\alpha}\right]\\ \simeq&\left[\Delta\mathcal{V}_{1}(f_{1})\right]^{\alpha}\rho\left[\left(\Delta\mathcal{V}_{1}(f_{1})\right)^{1-\alpha}\right]\end{split}$ (B.20)
for any $f_{1}$, where $c>0$ is a constant and $\rho(t)=\frac{1}{2}\left[(1+t)\log(1+t)+(1-t)\log(1-t)\right]$ for the logistic loss $\phi$.
Finally, combine Lemma LABEL:lem:surrogate.risk with (B.19), (B.20) and we get that
$\begin{split}&\left[\Delta\mathcal{V}_{1}(\widehat{f}_{1n})\right]^{\alpha}\rho\left[\left(\Delta\mathcal{V}_{1}(\widehat{f}_{1n})\right)^{1-\alpha}\right]+\mu_{1n}[\mathbb{E}{\mathds{1}}\{\widehat{f}_{1n}f_{2}^{*}<0\}-\mathbb{E}{\mathds{1}}\{f_{1}^{*}f_{2}^{*}<0\}-\widetilde{\delta}_{2N_{2}}(\tau)]\\ \lesssim&\mathbb{E}\ell_{1}\circ\widehat{f}_{1n}-\mathbb{E}\ell_{1}\circ f_{1}^{*}+\mathbb{E}\ell_{2}\circ\widehat{f}_{1n}-\mathbb{E}\ell_{2}\circ f_{1}^{*}\lesssim\delta_{1n}(\tau)\end{split}$
with probability at least $1-2e^{\tau}$, that is,
$\begin{split}\left[\Delta\mathcal{V}_{1}(\widehat{f}_{1n})\right]^{\alpha}\rho\left[\left(\Delta\mathcal{V}_{1}(\widehat{f}_{1n})\right)^{1-\alpha}\right]+\mu_{1n}[\mathbb{E}{\mathds{1}}\{\widehat{f}_{1n}f_{2}^{*}<0\}-\mathbb{E}{\mathds{1}}\{f_{1}^{*}f_{2}^{*}<0\}]\lesssim\delta_{1n}(\tau)+\mu_{1n}\widetilde{\delta}_{2N_{2}}(\tau).\end{split}$ (B.21)
By Taylor’s expansion, it is easy to see that $\rho(t)\simeq t^{2}$. Since $\mathbb{E}{\mathds{1}}\{f_{1}^{*}f_{2}^{*}<0\}\leq 1$, we can conclude that
$\left[\Delta\mathcal{V}_{1}(f_{1})\right]^{2-\alpha}\lesssim\delta_{1n}(\tau)+\mu_{1n}\widetilde{\delta}_{2N_{2}}(\tau)+\mu_{1n}$ (B.22)
and thus
$\Delta\mathcal{V}_{1}(f_{1})\lesssim(\delta_{1n}(\tau)+\mu_{1n})^{\frac{1}{2-\alpha}}$
with probability at least $1-2e^{\tau}$. ∎
B.6 Proof of Corollary B.2
Proof.
According to Lemma 5 and (9) of Bartlett et al., (2006),
$\mathbb{E}{\mathds{1}}\{f_{1}f_{1}^{*}<0\}\leq c\left[\mathbb{E}({\mathds{1}}\{f_{1}f_{1}^{*}<0\}\left\lvert 2\eta(\bm{X})-1\right\rvert)\right]^{\alpha}$
where $c$ is some constant. By Assumption B.5, we have
$2\eta(\bm{X})-1:=\frac{r_{1}^{(1)}(\bm{X})-r_{1}^{(-1)}(\bm{X})}{r_{1}^{(1)}(\bm{X})+r_{1}^{(-1)}(\bm{X})}\leq\frac{1}{c_{r}}\left\lvert r_{1}^{(1)}(\bm{X})-r_{1}^{(-1)}(\bm{X})\right\rvert.$
Since
$\begin{split}\Delta\mathcal{V}_{1}(f_{1})=&\mathbb{E}\left[\frac{R_{1}}{\pi(A;\bm{X})}{\mathds{1}}\{Af_{1}(\bm{X})<0\}\right]-\mathbb{E}\left[\frac{R_{1}}{\pi(A;\bm{X})}{\mathds{1}}\{Af_{1}^{*}(\bm{X})<0\}\right]\\ =&\mathbb{E}\left\{{\mathds{1}}\{f_{1}f_{1}^{*}<0\}\left\lvert r_{1}^{(1)}(\bm{X})-r_{1}^{(-1)}(\bm{X})\right\rvert\right\},\end{split}$
we can bound the disagreement rate by the value difference such that
$\mathbb{E}{\mathds{1}}\{f_{1}f_{1}^{*}<0\}\leq c\left[\frac{1}{c_{r}}\mathbb{E}({\mathds{1}}\{f_{1}f_{1}^{*}<0\}\left\lvert r_{1}^{(1)}(\bm{X})-r_{1}^{(-1)}(\bm{X})\right\rvert)\right]^{\alpha}\lesssim\left[\Delta\mathcal{V}_{1}(f_{1})\right]^{\alpha},$ (B.23)
that is, $\widehat{U}_{1n}\lesssim\left[\Delta\mathcal{V}_{1}(\widehat{f}_{1n})\right]^{\alpha}$. Then following (B.22) we have
$\widehat{U}_{1n}^{\frac{2-\alpha}{\alpha}}\lesssim\delta_{1n}(\tau)+\mu_{1n}\widetilde{\delta}_{2N_{2}}(\tau)+\mu_{1n}$
with probability at least $1-2e^{\tau}$ if we take $\rho(t)\simeq t^{2}$. ∎
B.7 Proof of Theorems LABEL:thm:convergence.rate.K and B.3
Proof.
The proof is similar to that for $K=2$. We only highlight the main differences here.
To extend the results of Lemma LABEL:lem:surrogate.risk to $K\geq 3$, we can write the surrogate loss for the fusion penalty as $\ell_{2}\circ f_{1}(\bm{Z}):=\mu_{1n}\sum_{k=2}^{K}\Omega_{1k}\psi_{\kappa_{1n}}[f_{1}(\bm{X})\widetilde{f}_{k}(\bm{X})]$. Then $\ell_{2}$ is Lipschitz continuous with respect to $f_{2}$ with Lipschitz constant $\frac{\mu_{1n}\sum_{k=2}^{K}\Omega_{1k}}{\kappa_{1n}}$. Hence, we have
$B:=2M\sqrt{M/\lambda_{1n}}+\mu_{1n}\sum_{k=2}^{K}\Omega_{1k}+M\simeq\lambda_{1n}^{-1/2}$
and
$\gamma_{n}:=\frac{B}{\sqrt{M/\lambda_{1n}}(M+\mu_{1n}\sum_{k=2}^{K}\Omega_{1k}\kappa_{1n}^{-1})}\simeq\frac{1}{1+\mu_{1n}\kappa_{1n}^{-1}},$
which shows that the conclusion in Lemma LABEL:lem:surrogate.risk still holds.
Now inequality (B.19) should be written as
$\begin{split}\mathbb{E}\ell_{2}\circ f_{1}-\mathbb{E}\ell_{2}\circ f_{1}^{*}=&\mu_{1n}\sum_{k=2}^{K}\Omega_{1k}\left[\mathbb{E}\psi_{\kappa_{n}}(f_{1}\widetilde{f}_{k})-\mathbb{E}\psi_{\kappa_{n}}(f_{1}^{*}\widetilde{f}_{k})\right]\\ \geq&\mu_{1n}\sum_{k=2}^{K}\Omega_{1k}\left[\mathbb{E}{\mathds{1}}\{f_{1}f_{k}^{*}<0\}-\mathbb{E}{\mathds{1}}\{f_{1}^{*}f_{k}^{*}<0\}-2\widetilde{\delta}_{kN_{k}}(\tau)\right],\end{split}$ (B.24)
with probability at least $1-(K-1)e^{\tau}$ for any $f_{1}$ by Assumption LABEL:asp:conv.secondary. Therefore, inequality (B.21) is changed to
$\begin{split}\left[\Delta\mathcal{V}_{1}(\widehat{f}_{1n})\right]^{\alpha}\rho\left[\left(\Delta\mathcal{V}_{1}(\widehat{f}_{1n})\right)^{1-\alpha}\right]+\mu_{1n}\sum_{k=2}^{K}[\mathbb{E}{\mathds{1}}\{\widehat{f}_{1n}f_{k}^{*}&<0\}-\mathbb{E}{\mathds{1}}\{f_{1}^{*}f_{k}^{*}<0\}]\\ &\lesssim\delta_{1n}(\tau)+\mu_{1n}\sum_{k=2}^{K}\widetilde{\delta}_{kN_{k}}(\tau)\end{split}$ (B.25)
with probability at least $1-Ke^{\tau}$, and thus
$\left[\Delta\mathcal{V}_{1}(f_{1})\right]^{2-\alpha}\lesssim\delta_{1n}(\tau)+\mu_{1n}\sum_{k=2}^{K}\widetilde{\delta}_{kN_{k}}(\tau)+\mu_{1n}.$ (B.26)
The inequality (B.3) follows from the assumption that $\sum_{k=2}^{K}\widetilde{\delta}_{kN_{k}}(\tau)=o(1)$.
Similarly, from (B.25) we can conclude that
$\mu_{1n}[\mathbb{E}{\mathds{1}}\{\widehat{f}_{1n}f_{k}^{*}<0\}-\mathbb{E}{\mathds{1}}\{f_{1}^{*}f_{k}^{*}<0\}]\lesssim\delta_{1n}(\tau)+\mu_{1n}\sum_{k=2}^{K}\widetilde{\delta}_{kN_{k}}(\tau)$
for any $k=2,\dots,K$ with probability at least $1-Ke^{\tau}$.
Combining results in (B.23) and (B.26), we obtain the bound
$\widehat{U}_{1n}^{\frac{2-\alpha}{\alpha}}\lesssim\delta_{1n}(\tau)+\mu_{1n}\sum_{k=2}^{K}\widetilde{\delta}_{kN_{k}}(\tau)+\mu_{1n}$
for the misclassification rate with probability at least $1-Ke^{\tau}$. ∎

### 09 Mar 2024 v3

@@
 In addition, we provide additional theoretical results and their proof in Section B.
-Appendix A ADDITIONAL DETAILS AND RESULTS FOR SECTIONS LABEL:sec:simulation AND LABEL:sec:realdata
-In this section, we provide some details about implementation and experiment results in the simulation study of Section LABEL:sec:simulation and in the real data analysis of Section LABEL:sec:realdata.
+Appendix A ADDITIONAL DETAILS AND EXPERIMENTS FOR SECTIONS LABEL:sec:simulation AND LABEL:sec:realdata
+In this section, we provide some details about implementation and experiments in the simulation study of Section LABEL:sec:simulation and the real data analysis of Section LABEL:sec:realdata.
 A.1 Implementation Details
@@
 The parameter $\sigma_{1n}$ in the Gaussian kernel is chosen as the median of the distances between all covariate pairs, which is a common heuristic when the sample size is not too large (Garreau et al.,, 2017).
-A.2 Additional Details for Data Generation and Value Function Estimation in Section LABEL:sec:simulation
+A.2 Details for Data Generation and Value Function Estimation in Section LABEL:sec:simulation
 The first two covariates are important variables that affect the outcomes and are generated as $X_{ij}\overset{i.i.d.}{\sim}Unif(-1,1)$ for all $i=1,\dots,n$ and $j=1,2$.
@@
 The misclassification rate of $\widehat{f}_{1n}$ is estimated by averaging ${\mathds{1}}\{\widehat{f}_{1n}f^{*}_{1}<0\}$ on the test set.
-A.3 Additional Results for Estimating FITR in Section LABEL:sec:simulation
+A.3 Additional Experiments for Estimating FITR in Section LABEL:sec:simulation
 For the scenarios considered in Section LABEL:sec:simulation, we present their true optimal values in Table A.1.
@@
 The coefficients are more similar for QIDS-change and CGI, which is expected since SAS measures the impacts of depression on social functioning.
-The Flanker variable has a different sign in the ITR of SAS than the other two variables.
+The Flanker variable has a different sign in the ITR of SAS than in the other two outcomes.
 Table A.4: Coefficients of the estimated ITRs of the three outcomes when linear kernel is used.
@@
 There exists $p_{0}>0$ such that $\pi(a;\bm{x})\equiv P(A=a|\bm{X}=\bm{x})>p_{0}$ for all $a\in\mathcal{A}$ and all $\bm{x}\in\mathcal{X}$.
-To bound the risk with 0-1 loss using the risk with convex surrogate loss, we need another common assumption in classification literature, which assumes the noise conditions around the decision boundary (Bartlett et al.,, 2006).
-For Assumption B.4, it has been shown that the boundary assumption of $\eta(\bm{x})$ regarding $\beta$ is equivalent to the misclassification assumption of $f_{1}$ regarding $\alpha$ (Bartlett et al.,, 2006).
+To bound the risk with 0-1 loss using the risk with convex surrogate loss, we need another common assumption in classification literature, the Tsybakov’s noise assumption, which assumes the noise conditions around the decision boundary (Bartlett et al.,, 2006).
 Assumption B.4.
@@
 for some $\beta>0$. Let $\alpha=\beta/(1+\beta)$ so that $\alpha\in(0,1]$.
+For Assumption B.4, it has been shown that the boundary assumption of $\eta(\bm{x})$ regarding $\beta$ is equivalent to the misclassification assumption of $f_{1}$ regarding $\alpha$ (Bartlett et al.,, 2006).
 Assumption B.5 is necessary when bounding the decision accuracy.
@@
 Under the weaker assumption, the convergence rate of $\mathbb{P}(\widehat{f}_{1n}f_{1}^{*}<0)$ is slower with a scaling constant $\alpha\in(0,1]$.
-B.2 Additional Details for Section LABEL:sec:theory
+B.2 Analysis Details for Section LABEL:sec:theory
 The parameters $\lambda_{1n},\mu_{1n},\kappa_{1n}$ should balance the estimation error and approximation error to achieve the minimal upper bound $\delta_{1n}(\tau)$.
-The minimal approximation error is $\left(\mu_{1n}\kappa_{1n}^{-1}\right)^{\frac{1}{1+q}}\lambda_{1n}^{\frac{q}{1+q}}$ when $\sigma_{1n}=\left(\mu_{1n}\kappa_{1n}^{-1}\lambda_{1n}^{-1}\right)^{\frac{1}{(1+q)d}}$.
-Then the best $\lambda_{1n}$ that balances the estimation error and the approximation error is $\lambda_{1n}=\left(\mu_{1n}\kappa_{1n}^{-1}\right)^{\frac{2(\Delta-1)+2(1+q)\nu}{3q+2\Delta+1}}n^{-\frac{1+q}{3q+2\Delta+1}}$.
+The minimal approximation error is $\gamma_{1n}^{\frac{1}{1+q}}\lambda_{1n}^{\frac{q}{1+q}}$ when $\sigma_{1n}=\left\{\gamma_{1n}\lambda_{1n}^{-1}\right\}^{\frac{1}{(1+q)d}}$.
+Then the best $\lambda_{1n}$ that balances the estimation error and the approximation error is $\lambda_{1n}=\gamma_{1n}^{\frac{2(\Delta-1)+2(1+q)\nu}{3q+2\Delta+1}}n^{-\frac{1+q}{3q+2\Delta+1}}$.
 When $\mu_{1n}=0$, the convergence rate of the risk is
@@
 $\begin{split}\mathbb{P}(f_{1}^{*}f_{2}^{*}>0)-\mathbb{P}(\widehat{f}_{1n}f_{2}^{*}>0)\leq\mathbb{P}(\widehat{f}_{1n}f_{1}^{*}<0)\lesssim n^{-\frac{\alpha}{2-\alpha}\frac{q}{3q+2\Delta+1}}.\end{split}$ (B.5)
+To compare the agreement rate with or without the fusion penalty, we break down equation (LABEL:equ:compare.SepL.FITR) as follows.
+When $\mu_{1n}\gtrsim\kappa_{1n}$, we have $\gamma_{1n}\simeq\mu_{1n}/\kappa_{1n}$.
+If $\alpha=1/2,\nu=3/2,\delta=1,\Delta=1/2$, the first condition in (LABEL:equ:compare.SepL.FITR) is satisfied with $\kappa_{1n}\gtrsim n^{-2q/(9q+6)},\mu_{1n}\gtrsim\kappa_{1n}$, and $N_{2}\gtrsim n^{3}\kappa_{1n}^{(9q+6)/q}\gtrsim n$.
+On the other hand, when $\mu_{1n}\lesssim\kappa_{1n}$, we have $\gamma_{1n}\simeq 1$.
+For the same $\alpha,\nu,\delta,\Delta$, the second condition in (LABEL:equ:compare.SepL.FITR) is satisfied with $\mu_{1n}\gtrsim n^{-2q/(9q+6)},\kappa_{1n}\gtrsim\mu_{1n}$, and $N_{2}\gtrsim n^{3}\mu_{1n}^{(9q+6)/q}\gtrsim n$.
 B.3 Proof of Lemma LABEL:lem:surrogate.risk
@@
 when $\lVert\widetilde{f}_{2}\rVert_{\infty}=1$. Combining (B.9), (B.10) and (B.11), we can bound the approximation error (B.8) as
-$\inf_{f_{1}\in\mathcal{H}}\left[\mathcal{R}(f_{1})+\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}\right]-\mathcal{R}(f_{1}^{*})\lesssim\lambda_{1n}\sigma_{1n}^{d}+\left(1+\frac{\mu_{1n}}{\kappa_{1n}}\right)(2d)^{qd/2}\sigma_{1n}^{-qd}.$ (B.12)
+$\inf_{f_{1}\in\mathcal{H}}\left[\mathcal{R}(f_{1})+\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}\right]-\mathcal{R}(f_{1}^{*})\lesssim\lambda_{1n}\sigma_{1n}^{d}+\gamma_{1n}(2d)^{qd/2}\sigma_{1n}^{-qd},$ (B.12)
+where $\gamma_{1n}:=1+\mu_{1n}/\kappa_{1n}$.
 To bound (B.7), we will use Talagrand’s inequality quoted as follows (Steinwart and Scovel,, 2007, Theorem 5.3).
@@
 $\lambda_{1n}\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}\leq\mathbb{P}_{n}(\ell_{1}\circ f_{1}+\ell_{2}\circ f_{1})\leq\frac{r}{p_{0}}+\mu_{1n}\Omega_{12}.$
-Since $r/p_{0}+\mu_{1n}\Omega_{12}\simeq M$, $\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}$ is bounded by $\sqrt{M/\lambda_{1n}}$, where $M:=r/p_{0}$.
+Since $r/p_{0}+\mu_{1n}\Omega_{12}\simeq M$ with $M:=r/p_{0}$, $\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}$ is bounded by $\sqrt{M/\lambda_{1n}}$.
 To this end, it suffices to consider the ball of radius $\sqrt{M/\lambda_{1n}}$.
@@
 With $u,u^{\prime}\in B^{-1}\left\{\ell_{1}\circ f_{1}+\ell_{2}\circ f_{1}:f\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})\right\}$,
-$\begin{split}&\log N\left(B^{-1}\left\{\ell_{1}\circ f_{1}+\ell_{2}\circ f_{1}:f\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})\right\},\epsilon,L_{2}(\mathbb{P}_{n})\right)\\ \leq&\log N\left(B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}}),\frac{B\epsilon}{M+\mu_{1n}\Omega_{12}\kappa_{1n}^{-1}},L_{2}(\mathbb{P}_{n})\right)\\ \leq&\log N\left(B_{\mathcal{H}},\gamma_{n}\epsilon,L_{2}(\mathbb{P}_{n})\right),\end{split}$
-where
-$\gamma_{n}:=\frac{B}{\sqrt{M/\lambda_{1n}}(M+\mu_{1n}\Omega_{12}\kappa_{1n}^{-1})}\simeq\frac{1}{1+\mu_{1n}\kappa_{1n}^{-1}}$
+$\begin{split}&\log N\left(B^{-1}\left\{\ell_{1}\circ f_{1}+\ell_{2}\circ f_{1}:f\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})\right\},\epsilon,L_{2}(\mathbb{P}_{n})\right)\\ \leq&\log N\left(B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}}),\frac{B\epsilon}{M+\mu_{1n}\Omega_{12}\kappa_{1n}^{-1}},L_{2}(\mathbb{P}_{n})\right)\\ \leq&\log N\left(B_{\mathcal{H}},\gamma_{1n}^{-1}\epsilon,L_{2}(\mathbb{P}_{n})\right),\end{split}$
+since
+$\frac{B}{\sqrt{M/\lambda_{1n}}(M+\mu_{1n}\Omega_{12}\kappa_{1n}^{-1})}\simeq\frac{1}{1+\mu_{1n}\kappa_{1n}^{-1}}=\gamma_{1n}^{-1}.$
 For the second term on the right-hand side, it follows that
@@
 since $\lambda_{1n}\lVert f_{1}\rVert_{\mathcal{H}}^{2}\leq M$ for all $f\in B_{\mathcal{H}}(\sqrt{M/\lambda_{1n}})$. Therefore, we can conclude that
-$\log N(\epsilon,B^{-1}\mathcal{G},L_{2}(\mathbb{P}_{n}))\leq\log N\left(B_{\mathcal{H}},\gamma_{n}\epsilon,L_{2}(\mathbb{P}_{n})\right)+\log\frac{M}{B\epsilon}.$
+$\log N(\epsilon,B^{-1}\mathcal{G},L_{2}(\mathbb{P}_{n}))\leq\log N\left(B_{\mathcal{H}},\gamma_{1n}^{-1}\epsilon,L_{2}(\mathbb{P}_{n})\right)+\log\frac{M}{B\epsilon}.$
 Theorem 2.1 of Steinwart and Scovel, (2007) then yields that
-$\sup_{\mathbb{P}_{n}}\log N(\epsilon,B^{-1}\mathcal{G},L_{2}(\mathbb{P}_{n}))\lesssim\sigma_{1n}^{(1-\nu/2)(1+\delta)d}(\gamma_{n}\epsilon)^{-\nu},$
+$\sup_{\mathbb{P}_{n}}\log N(\epsilon,B^{-1}\mathcal{G},L_{2}(\mathbb{P}_{n}))\lesssim\sigma_{1n}^{(1-\nu/2)(1+\delta)d}(\gamma_{1n}^{-1}\epsilon)^{-\nu},$
 where $\sigma_{1n}>0$ is the parameter of the Gaussian kernel associated with $\mathcal{H}$, and $00<\nu\leq 2,\delta>0,\epsilon>$ 0.
-Therefore, we have $a=\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{n}^{-\nu}$, $p=\nu$ and
+Therefore, we have $a=\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{1n}^{\nu}$, $p=\nu$ and
 $\operatorname{Rad}(\mathcal{G},n,\epsilon)\leq c_{p}\max\left\{B^{\frac{p}{2}}\epsilon^{\frac{1}{2}-\frac{p}{4}}\left(\frac{a}{n}\right)^{\frac{1}{2}},B\left(\frac{a}{n}\right)^{\frac{2}{2+p}}\right\}.$
 With $\epsilon=4B^{2}$, we can bound the modulus of continuity as
-$\omega_{n}(\mathcal{G},\epsilon)\leq 2\operatorname{Rad}(\mathcal{G},n,\epsilon)\lesssim Ba^{\frac{1}{2}}n^{-\frac{1}{2}}\simeq\lambda_{1n}^{-\frac{1}{2}}\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{n}^{-\nu}n^{-\frac{1}{2}}.$ (B.15)
+$\omega_{n}(\mathcal{G},\epsilon)\leq 2\operatorname{Rad}(\mathcal{G},n,\epsilon)\lesssim Ba^{\frac{1}{2}}n^{-\frac{1}{2}}\simeq\lambda_{1n}^{-\frac{1}{2}}\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{1n}^{\nu}n^{-\frac{1}{2}}.$ (B.15)
 The definition of $\widehat{f}_{1n}$ yields that
@@
 Plugging (B.15) into (B.14) and we have
-$\begin{split}&\left[\mathcal{R}(\widehat{f}_{1n})+\lambda_{1n}\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}\right]-\left[\mathcal{R}(f_{1}^{\dagger})+\lambda_{1n}\lVert f_{1}^{\dagger}\rVert_{\mathcal{H}}^{2}\right]\leq\sup_{h\in\mathcal{E}}\mathbb{P}_{n}h\\ \lesssim&\lambda_{1n}^{-\frac{1}{2}}\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{n}^{-\nu}n^{-\frac{1}{2}}+\sqrt{\frac{2\tau\lambda_{1n}^{-1}}{n}}+\frac{\tau\lambda_{1n}^{-\frac{1}{2}}}{n}\\ \lesssim&\lambda_{1n}^{-\frac{1}{2}}n^{-\frac{1}{2}}\left[\sqrt{\tau}+\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{n}^{-\nu}\right]\end{split}$ (B.16)
+$\begin{split}&\left[\mathcal{R}(\widehat{f}_{1n})+\lambda_{1n}\lVert\widehat{f}_{1n}\rVert_{\mathcal{H}}^{2}\right]-\left[\mathcal{R}(f_{1}^{\dagger})+\lambda_{1n}\lVert f_{1}^{\dagger}\rVert_{\mathcal{H}}^{2}\right]\leq\sup_{h\in\mathcal{E}}\mathbb{P}_{n}h\\ \lesssim&\lambda_{1n}^{-\frac{1}{2}}\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{1n}^{\nu}n^{-\frac{1}{2}}+\sqrt{\frac{2\tau\lambda_{1n}^{-1}}{n}}+\frac{\tau\lambda_{1n}^{-\frac{1}{2}}}{n}\\ \lesssim&\lambda_{1n}^{-\frac{1}{2}}n^{-\frac{1}{2}}\left[\sqrt{\tau}+\sigma_{1n}^{(1-\nu/2)(1+\delta)d}\gamma_{1n}^{\nu}\right]\end{split}$ (B.16)
 with probability at least $1-e^{-\tau}$ for any $\tau\geq 1$.
@@
 and
-$\gamma_{n}:=\frac{B}{\sqrt{M/\lambda_{1n}}(M+\mu_{1n}\sum_{k=2}^{K}\Omega_{1k}\kappa_{1n}^{-1})}\simeq\frac{1}{1+\mu_{1n}\kappa_{1n}^{-1}},$
+$\gamma_{1n}^{-1}:=\frac{B}{\sqrt{M/\lambda_{1n}}(M+\mu_{1n}\sum_{k=2}^{K}\Omega_{1k}\kappa_{1n}^{-1})}\simeq\frac{1}{1+\mu_{1n}\kappa_{1n}^{-1}},$
 which shows that the conclusion in Lemma LABEL:lem:surrogate.risk still holds.